# sharp bottom spectrum and scalar curvature rigidity

### v1

@@ Abstract.
We prove a sharp upper bound for the bottom spectrum of Laplacian on geometrically contractible manifolds with scalar curvature lower bound, and characterize the distribution of scalar curvature when equality holds. Moreover, we prove a scalar curvature rigidity theorem if the manifold is the universal cover of a closed hyperbolic manifold.
The first author is partially supported by NSF 1952693.
@@ 1. Introduction
Let $(X^{n},g)$ be a connected, complete, noncompact Riemannian manifold, and $\Delta$ the corresponding Beltrami Laplacian on $(X^{n},g)$ defined by
@@
The $L^{2}$-bottom spectrum of $\Delta$ is defined by
$\lambda_{1}(X,g)=\inf\left\{\frac{\int_{X}|\nabla f|^{2}}{\int_{X}f^{2}}:\ f\in C_{c}^{\infty}(M),\ f\neq 0\right\}$ (1.1)
Recall that the classical comparison theorem in [Cheng_eigenvalue] implies that, if $(X^{n},g)$ satisfies $\mathrm{Ric}_{g}\geq-(n-1)$, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}.$
A natural question is to generalize this question to scalar curvature lower bound. Recently, Munteanu–Wang in [MunteanuWang, Theorem 1.1] generalize the sharp bottom spectrum estimate on three-dimensional manifolds with scalar curvature lower bound. If $(X^{3},g)$ is a complete Riemannian manifold with scalar curvature $\mathrm{Sc}_{g}\geq-6$ and satisfies either one of the following properties:
@@
In this paper, we further study the bottom spectrum under scalar curvature lower bound using the Dirac operator method and the higher index theory. Higher index theory [BaumConnesHigson, MR0996437, Roe] is a far-reaching generalization of the classical Fredholm index, particularly for non-compact manifolds, and is a more refined index theory than the theorems of Atiyah–Singer. We will later give a brief introduction for the basic setups for higher indices in Section 2.
Let us recall that a Riemannian manifold is said to be geometrically contractible if there exists a function $R(r)\geq r$ for any $r\geq 0$ such that $B(p,r)$ is contractible in $B(p,R(r))$ for any $p\in M$. Note that the universal cover of any aspherical manifold is geometrically contractible (see [wang2024fillingradiusquantitativektheory, Example 2.6]). A complete Riemannian manifold is said to have bounded geometry if the sectional curvature and its derivatives are uniformly bounded and injective radius has uniformly lower bound, i.e., $|\nabla^{\alpha}\sec_{g}|\leq K_{\alpha},\ \operatorname{Inj}(M)\geq i>0$ for any multi-index $\alpha$ and constants $K_{\alpha},i>0$. The first main theorem of this paper is as follows.
@@
Suppose that $(X^{n},g)$ is complete, geometrically contractible Riemannian manifold with bounded geometry and scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)\kappa$ for some constant $\kappa\geq 0$. If $(X^{n},g)$ satisfies the coarse Novikov conjecture, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}\kappa.$ (1.2)
@@
The precise statement of the coarse Novikov conjecture is in Conjecture 2.6. Recall that a subset $S$ in $X$ is is said to be a net of $X$ if there exists $r>0$ such that $N_{r}(S)=X$, where $N_{r}(S)=\{x\in X:\mathrm{dist}(x,S)<r\}$. Theorem 1.1 is a geometric version of a more general theorem presented in Section 4. We emphasize that the upper bound $\frac{(n-1)^{2}}{4}\kappa$ is sharp, as the standard hyperbolic space $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$ has scalar curvature $-n(n-1)$ and $\lambda_{1}(\mathbb{H}^{n},g_{\mathbb{H}^{n}})=\frac{(n-1)^{2}}{4}$.
Classically, Li-Wang make use of the harmonic function theory to study the sharpness of the upper bound on the bottom spectrum and the splitting rigidity in [Li_Wang_positive_spectrum_1, Li_Wang_positive_spectrum_2] if the Ricci curvature $\mathrm{Ric}_{g}\geq-(n-1)$. As pointed out in [MunteanuWang], it is impossible to generalize the classical sharp bottom spectrum result to any complete Riemannian manifold with only scalar curvature assumption. Essentially, our philosophy for proving Theorem 1.1 is based on the Novikov conjecture. Moreover, the method, using higher index theory as a starting point, has a broader application since it is easier to achieve non-vanishing higher index than non-vanishing Fredholm index or $L^{2}$-index used in [Davaux] and has no restriction on the parity of the dimension. The rigidity result [Davaux, Theorem C] is modeled on the flat space due to the assumption of nonzero $\hat{A}$-genus. See the survey related to Novikov conjecture in [Xie_Yu_higher_invariants] or a list at the end of the Section 2 in this paper.
Furthermore, the net characterization on general complete noncompact manifold, as equality holds, is possibly the best expectation and cannot be further improved in general, which is supported by the following example inspired by [MunteanuWang, Theorem 1.4 & Example 1.5].
@@
Moreover, if the Riemannian manifold is the universal cover of a closed manifold, we further obtain a scalar curvature rigidity. Here, we only state the theorem in case of closed hyperbolic manifold.
@@
Suppose that $M^{n}$ is a closed hyperbolic manifold. If $(M^{n},g)$ has scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)\kappa$ for some constant $\kappa\geq 0$, then
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{(n-1)^{2}}{4}\kappa,$ (1.3)
where $(\widetilde{M},\widetilde{g})$ is the Riemannian universal cover of $(M,g)$. Moreover, if the equality holds, then $Sc_{g}=-n(n-1)\kappa$.
A more general result is given Section 4, showing that the conclusion also holds for closed essential manifolds whose fundamental groups satisfy the strong Novikov conjecture 2.7.
It is known that closed hyperbolic manifold admits no complete Riemannian metric with non-negative scalar curvature. Note that Theorem 1.3 can be rewritten as
$\inf_{p\in M}(-\mathrm{Sc}_{g}(p))\leq\frac{4n}{n-1}\lambda_{1}(\widetilde{M},\widetilde{g}).$
Hence, it can be viewed as a quantitative version of the nonexistence of complete Riemannian metric with non-negative scalar curvature on hyperbolic manifold. A further question in this direction is to find an optimal upper bound on $\inf_{p\in M}(-\mathrm{Sc}_{g}(p))$ in terms of the group complexity invariant of the fundamental group for hyperbolic manifold or aspherical manifold, and the $\operatorname{\mathrm{vol}}(M)$. Moreover, it is interesting to study the relations between the bottom spectrum $\lambda_{1}$ and the macroscopic scalar curvature, the seminal works [Guth_volumes_of_balls, Roman_volume, B_K_macroscopic, Sabine_Roman_volume, Hannah_Kei_Area, alpert2024growthuniversalcoverlarge] provide series of relations between the macroscopic curvature and other geometric and topological invariants. Finally, we note that Wang [Wang_eigenvalue] prove that, if $\mathrm{Ric}_{g}\geq-(n-1)$, the $\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{(n-1)^{2}}{4}$. In particular, the equality holds if and only if $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$. Theorem 1.3 directly leads to a further rigidity conjecture as follows.
@@
then $(\widetilde{M},\widetilde{g})$ isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$.
Note that Conjecture 1.4 has been confirmed in [MunteanuWang] for $n=3$. Finally, Theorem 1.3 is connected with other problems as follows. It is known that for closed Riemannian manifold,
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{h^{2}_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})}{4}$
with $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})$ the volume entropy. Ledrappier–Wang and Liu in [Fran_Wang_entropy_rigidity, Liu_volume_entropy] prove that, if $(M^{n},g)$ is a closed Riemannian manifold with $\mathrm{Ric}_{g}\geq-(n-1)$, then $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\leq n-1$, and the equality holds if and only if $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$. However, Kazaras–Song–Xu in [kazaras2024scalarcurvaturevolumeentropy] prove that the volume entropy $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\leq n-1$ does not hold for a closed hyperbolic manifold $(M^{n},g)$ with scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)$. Additionally, it’s known that there exists a constant $c_{n}$ such that
$\big{(}h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\big{)}^{n}\operatorname{\mathrm{vol}}(M)\geq c_{n}\|M\|$
for any closed Riemannian manifold by Gromov in [Gromov_volume], and also a sharp constant $c_{n}$ has been obtained for the locally symmetric spaces due to Besson–Courtois–Gallot in [BCG_entropy, BCG_minimal_entropy]. Motivated by this, we propose the following question.
Problem 1.5.
@@
  • We expect that there exists a constant $c_{n}$ such that
@@
Problem 1.5 holds trivially when $\pi_{1}(M)$ is amenable, where $\|M\|=\lambda_{1}(\widetilde{M},\widetilde{g})=0$. Moreover, Mohsen studies the first question in [Olivier_bottom_spectrum_hyperbolic, Theorem 1] and proves that the hyperbolic metric on $M$ is a saddle point of $\lambda_{1}(\widetilde{M})$. More precisely, the bottom spectrum is maximal among the conformal metrics of same volume, and minimal in its Ebin class. Problem 1.5 is connected with Gromov simplicial volume conjecture for general closed Riemannian manifold (see [Gromov_101, Section 26]) and Schoen conjecture on closed hyperbolic manifold (see [Schoen_variational_theory]), and macroscopic scalar curvature is a stronger concept than scalar curvature and many progresses has been made(see [Guth_volumes_of_balls, alpert2024growthuniversalcoverlarge, Sabine_Roman_volume] and the literature therein) in this direction. Hence, it is worthwhile of studying Problem 1.5 for its own interest and applications.
@@
This paper is structured as follows: In Section 2, we provide the necessary background on higher index theory. In Section 3, we establish the Kato inequality for harmonic spinors, which is a standard result. We give a elementary proof and a slightly more general version for non-compact manifolds. In Section 4, we prove the main theorems presented in this paper and derive several related corollaries. Finally, in Section 5, to establish a net characterization for complete noncompact manifolds and demonstrate scalar curvature rigidity for cocompact manifolds, we prove a unique continuation theorem on complete Riemannian manifolds.
@@ 2. Preliminaries on higher index theory
In this section, we review the construction of the geometric $C^{*}$-algebras and the higher index. The readers may refer to [willett2020higher] for more details.
@@ 3. Kato inequality for harmonic spinor
Let $(X^{n},g)$ be a complete Riemannian manifold. Let $E$ be a vector bundle over $X$ equipped with a Clifford action of $TX$. Let $D$ be the Dirac operator
@@
Let $V$ be a vector space. Let $W$ be a vector space equipped with a $Cl(V)$-action. Set
$T\colon V\otimes W\to W,~{}\psi\mapsto\sum_{i=1}^{n}c(e_{i})\langle\psi,e_{i}\rangle.$
Then for any $\psi\in\ker T$ and $\xi\in V$, we have
@@
Let $E$ be a bundle over $X^{n}$ equipped with a Clifford action of $TX$. There exists $c_{n}>0$ that only depends on $n$ such that for any smooth section $\xi$ of $E$, we have

### v2

@@ Abstract.
We prove a sharp upper bound on the bottom spectrum of Beltrami Laplacian on geometrically contractible Riemannian manifolds with scalar curvature lower bound, and then characterize the distribution of the scalar curvature when the equality holds. Moreover, we prove a scalar curvature rigidity theorem if the manifold is the universal cover of a closed hyperbolic manifold.
@@ 1. Introduction
Suppose that $(X^{n},g)$ is a connected, complete, noncompact Riemannian manifold and $\Delta$ is the corresponding Beltrami Laplacian on $(X^{n},g)$ defined as
@@
The $L^{2}$-bottom spectrum of $\Delta$ on $(X,g)$ is defined by (see [Cheng_eigenvalue]*Section 4 or [Li_geometric_analysis]*Definition 6.3)
$\lambda_{1}(X,g)=\inf\left\{\frac{\int_{X}|\nabla f|^{2}}{\int_{X}f^{2}}:\ f\in C_{c}^{\infty}(M),\ f\neq 0\right\}.$ (1.1)
Recall that Cheng, using the classical comparison theorem (see [Cheng_eigenvalue]*Theorem 4.2) proves that, if $(X^{n},g)$ satisfies $\mathrm{Ric}_{g}\geq-(n-1)$, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}.$ (1.2)
A further question is to generalize estimate (1.2) on manifolds with scalar curvature lower bound. In this direction, using the harmonic function theory, Munteanu–Wang generalize the sharp bottom spectrum estimate for three-dimensional manifolds with (negative) scalar curvature lower bound as follows.
Theorem 1.1 (Munteanu-Wang, see [MunteanuWang]*Theorem 1.1).
Suppose that $(X^{3},g)$ is a complete, noncompact, three dimensional Riemannian manifold with scalar curvature $\mathrm{Sc}_{g}\geq-6$. If $X$ satisfies either one of the following properties:
@@
In this paper, we further study and then extend the sharp bottom spectrum estimate to complete Riemannian manifold of higher dimension using the Dirac operator and higher index theory. In our context, the Dirac operator and higher index theory can be (technically) viewed as a parallel of the Laplacian operator and the harmonic function theory on complete noncompact Riemannian manifold that as in [MunteanuWang].
Recall that Li-Wang initiated to make use of the harmonic function theory to study the sharpness of the upper bound on the bottom spectrum of Beltrami-Laplacian and the splitting rigidity on complete, noncompact Riemannian manifold with Ricci curvature $\mathrm{Ric}_{g}\geq-(n-1)$ (see [Li_Wang_positive_spectrum_1, Li_Wang_positive_spectrum_2]). As it is pointed out in [MunteanuWang], Theorem 1.1 does not hold without any topological or geometric assumption (see [MunteanuWang]*Example 1.2). Hence, finding a sufficient topological or geometric condition is a necessary step to generalize the classical sharp bottom spectrum estimate on complete Riemannian manifolds with only scalar curvature constraint.
Let us recall that a complete Riemannian manifold $(X^{n},g)$ is said to be geometrically contractible if there exists a function $R(r)\geq r$ for any $r\geq 0$ such that $B(p,r)$ is contractible in $B(p,R(r))$ for any $p\in X$. Note that the universal Riemannian cover of any closed, aspherical Riemannian manifold is geometrically contractible (see [wang2024fillingradiusquantitativektheory]*Example 2.6). Moreover, a complete Riemannian manifold $(X^{n}g)$ is said to be bounded geometry if the sectional curvature and its derivatives are uniformly bounded, and the injective radius has uniformly lower bound, i.e., $|\nabla^{\alpha}\sec_{g}|\leq K_{\alpha},\ \operatorname{Inj}(M)\geq i>0$ for any multi-index $\alpha$ and constants $K_{\alpha},i>0$.
Now we are ready to state the first main theorem in this paper as follows.
@@
Suppose that $(X^{n},g)$ is a complete, geometrically contractible Riemannian manifold with bounded geometry and scalar curvature $\mathrm{Sc}_{g}\geq-\kappa$ for some constant $\kappa\geq 0$. If $(X^{n},g)$ satisfies the coarse Novikov conjecture, then
$\lambda_{1}(X,g)\leq\frac{n-1}{4n}\kappa.$ (1.3)
@@
See Conjecture 2.6 in Section 2 for the precise statement of the coarse Novikov conjecture and more general version of Theorem 1.2 in Section 4. Recall that a subset $S$ in $X$ is is said to be a net of $X$ if there exists $r>0$ such that $N_{r}(S)=X$, where $N_{r}(S)=\{x\in X:\mathrm{dist}(x,S)<r\}$. Theorem 1.2 is a geometric version of a more general theorem presented in Section 4. We emphasize that the upper bound $\frac{n-1}{4n}\kappa$ is sharp since the standard hyperbolic space $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$ has scalar curvature $-n(n-1)$ and $\lambda_{1}(\mathbb{H}^{n},g_{\mathbb{H}^{n}})=\frac{(n-1)^{2}}{4}$.
Here, our philosophy of approaching Theorem 1.2 is grounded in the topological condition: (coarse) Novikov conjecture. Moreover, on one hand, the method of using the higher index theory is parallel to using the harmonic function theory developed in [Li_Wang_positive_spectrum_1, Li_Wang_positive_spectrum_2, MunteanuWang], on the other hand, in comparison with the classical index theory, the higher index theory has a broader application since it is easier to achieve non-vanishing higher index than non-vanishing Fredholm index or $L^{2}$-index used in [Davaux] and has no restriction on the parity of the dimension in [MunteanuWang]. Note that the rigidity theorem (see [Davaux, Theorem C]) is modeled on the flat Riemannian manifold due to the assumption of nonzero $\hat{A}$-genus.
Moreover, the net characterization on general complete, noncompact Riemannian manifold, as equality holds, is possibly the best expectation and cannot be further improved in general, which is supported by the following example inspired by [MunteanuWang, Theorem 1.4 & Example 1.5].
@@
Furthermore, if the Riemannian manifold is the universal cover of a closed manifold, we further obtain a scalar curvature rigidity. Here, we only state the theorem in case of closed hyperbolic manifold (see more general version in Section 4).
@@
Suppose that $M^{n}$ is a closed hyperbolic manifold. If the complete Riemannian manifold $(M^{n},g)$ has scalar curvature $\mathrm{Sc}_{g}\geq-\kappa$ for some constant $\kappa\geq 0$, then
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{n-1}{4n}\kappa,$ (1.4)
where $(\widetilde{M},\widetilde{g})$ is the Riemannian universal cover of $(M,g)$. Moreover, if the equality holds, then $\mathrm{Sc}_{g}=-\kappa$.
Recall that any closed hyperbolic manifold admits no complete Riemannian metric with nonnegative scalar curvature (see [Gromov_four_lectures, Section 4.1.2]). Note that Theorem 1.4 can be rewritten as
$\inf_{p\in M}(-\mathrm{Sc}_{g}(p))\leq\frac{4n}{n-1}\lambda_{1}(\widetilde{M},\widetilde{g}).$ (1.5)
Hence, the inequality (1.5) can be viewed as a quantitative version of nonexistence of complete Riemannian metric with nonnegative scalar curvature on a closed hyperbolic manifold. A further question in this direction is to find an optimal upper bound on $\inf_{p\in M}(-\mathrm{Sc}_{g}(p))$ in terms of the group complexity invariant of the fundamental group for hyperbolic manifolds or aspherical manifolds, and the $\operatorname{\mathrm{vol}}(M)$. Also, it is interesting to study the relationship between the bottom spectrum $\lambda_{1}$ and the macroscopic scalar curvature, the seminal works [Guth_volumes_of_balls, Roman_volume, B_K_macroscopic, Sabine_Roman_volume, Hannah_Kei_Area, alpert2024growthuniversalcoverlarge] provide series of relations between the macroscopic curvature and other geometric and topological invariants. Moreover, X.D. Wang proves that, if $\mathrm{Ric}_{g}\geq-(n-1)$, then
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{(n-1)^{2}}{4}.$
In particular, the equality holds if and only if $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$ (see [Wang_eigenvalue]*Theorem 1.4). Hence, Theorem 1.4 directly leads us a further rigidity conjecture as follows.
@@
then $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$.
Note that Munteanu–Wang confirm Conjecture 1.5 for $n=3$ (see [MunteanuWang]*Theorem 1.3).
Finally, Theorem 1.4 is connected with other problems as follows. Recall that, for closed Riemannian manifold, we have (see [Brooks_spectrum_volume]*Theorem 1)
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{h^{2}_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})}{4}.$
with $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})$ the volume entropy defined by
$h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})=\lim_{R\to\infty}\frac{\log(\widetilde{B}(\tilde{p},R))}{R}.$ (1.6)
Here, $\widetilde{B}(\tilde{p},R)$ is the geodesic ball in $(\widetilde{M},\widetilde{g})$ with center $\tilde{p}\in\widetilde{M}$ and radius $R$. Ledrappier–Wang prove that, if $(M^{n},g)$ is a closed Riemannian manifold with $\mathrm{Ric}_{g}\geq-(n-1)$, then $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\leq n-1$, and the equality holds if and only if $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$ (see [Fran_Wang_entropy_rigidity, Liu_volume_entropy]). However, Kazaras–Song–Xu prove that the volume entropy $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\leq n-1$ does not hold for closed hyperbolic manifold $(M^{n},g)$ with scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)$ (see [kazaras2024scalarcurvaturevolumeentropy]*Theorem 0.2). Moreover, Gromov proves that, for any closed Riemannian manifold $(M,g)$ (see [Gromov_volume]), there exists a constant $c_{n}>0$ such that
$\big{(}h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\big{)}^{n}\operatorname{\mathrm{vol}}(M)\geq c_{n}\|M\|,$
and a sharp constant $c_{n}$ has been proved for the locally symmetric spaces by Besson–Courtois–Gallot (see [BCG_entropy, BCG_minimal_entropy]). Motivated by these connections, we propose a problem as follows.
Problem 1.6.
@@
  • We conjecture that there exists a constant $c_{n}$ such that
@@
Note that, if $\pi_{1}(M)$ is amenable, then Problem 1.6 holds. Moreover, Mohsen studies the first question in [Olivier_bottom_spectrum_hyperbolic, Theorem 1] and proves that the hyperbolic metric on $M$ is a saddle point of $\lambda_{1}(\widetilde{M})$. More precisely, the bottom spectrum is maximal among the conformal metrics of same volume, and minimal in its Ebin class. Problem 1.6 is connected with Gromov simplicial volume conjecture for general closed Riemannian manifold (see [Gromov_101, Section 26]) and Schoen conjecture on closed hyperbolic manifold (see [Schoen_variational_theory]). Note that macroscopic scalar curvature is a stronger concept than scalar curvature and many progresses has been made(see [Guth_volumes_of_balls, alpert2024growthuniversalcoverlarge, Sabine_Roman_volume] and the literature therein) in this direction. Hence, it is worthwhile of studying Problem 1.6 for its own interest and applications.
@@
This paper is structured as follows: In Section 2, we provide the necessary background on higher index theory. In Section 3, we establish the Kato inequality for harmonic spinors, which is a standard result. We give an elementary proof and a slightly more general version for noncompact manifolds. In Section 4, we prove the main theorems presented in this paper and derive several related corollaries. Finally, in Section 5, to establish a net characterization for complete noncompact manifolds and demonstrate scalar curvature rigidity for cocompact manifolds, we prove a unique continuation theorem on complete Riemannian manifolds.
@@ 2. Preliminaries on higher index theory
In this section, we will review the construction of the geometric $C^{*}$-algebras and the higher index (see the textbook [willett2020higher] for more details). The higher index theory [BaumConnesHigson, MR0996437, Roe] is a far-reaching generalization of the classical Fredholm index, particularly for non-compact manifolds, and is a more refined index theory than the theorems of Atiyah–Singer.
@@ 3. Kato inequality for harmonic spinor
Suppose that $(X^{n},g)$ is a complete Riemannian manifold and $E$ is a vector bundle over $X$ equipped with a Clifford action of $TX$. Let $D$ be the Dirac operator
@@
Suppose that $V$ is a vector space and $W$ is a vector space equipped with a $Cl(V)$-action. Let
$T\colon V\otimes W\to W,~{}\psi\mapsto\sum_{i=1}^{n}c(e_{i})\langle\psi,e_{i}\rangle,$
then, for any $\psi\in\ker T$ and $\xi\in V$, we have
@@
If $E$ is a bundle over $X^{n}$ equipped with a Clifford action of $TX$, then there exists $c_{n}>0$ depending only on $n$ such that, for any smooth section $\xi$ of $E$, we have